A weather balloon is inflated to a volume of 28.6 L at a pressure of 741 mmHg and a temperature of 28.5 ∘C. The balloon rises in the atmosphere to an altitude where the pressure is 400. mmHg and the temperature is -14.9 ∘C.
Answer #1

Given:

Pi = 741 mmHg

Pf = 400 mmHg

Vi = 28.6 L

Ti = 28.5 oC

= (28.5+273) K

= 301.5 K

Tf = -14.9 oC

= (-14.9+273) K

= 258.1 K

use:

(Pi*Vi)/(Ti) = (Pf*Vf)/(Tf)

(741 mmHg*28.6 L)/(301.5 K) = (400 mmHg*Vf)/(258.1 K)

Vf = 45.4 L

Answer: 45.4 L